Hello..... A body cools from 80 C to 50 C in 5 min. Calculate the time it takes to cool from 60 C to 30 C . The temperature of the surroundings is 20 C.

Hello.....
       A body cools from 80 C to 50 C in 5 min. Calculate the time it takes to cool from 60 C to 30 C . The temperature of the surroundings is 20 C.

Grade:9

1 Answers

Vikas TU
7 years ago
From
-dT/dt = k(T – T0)
dT = – k(T – T0)dt
Integerating both sides with appropriate limits,
log(T- T0) = -Kt
(-2.306/K)log2t = K = -2.303*log2/300
we get
t = 10 minutes.